OpenStudy (idealist10):

Find the value of B: -12Axe^(-x)+2Ae^(-x)-6Be^(-x)=-6xe^(-x)

OpenStudy (idealist10):

I know that A=1/2 but what's B?

OpenStudy (idealist10):

@ganeshie8 @mathstudent55 @ParthKohli @texaschic101

OpenStudy (idealist10):

@phi @Hero @Directrix

OpenStudy (phi):

you should factor out a 2 e^(-x) from each term on both sides and then cancel or if you like, multiply both sides by 1/2 e^x what do you get ?

OpenStudy (phi):

after dividing by 2e^(-x) -12Axe^(-x)+2Ae^(-x)-6Be^(-x)=-6xe^(-x) you get -6Ax +A -3B= -3x now match up corresponding terms -6Ax = -3 x so A = 1/2 also -3B+A = 0 so you can solve for B.

OpenStudy (idealist10):

Thank you!